Write the equation of a line through the points (4,3) and (6,4). Show work.
Answer by stanbon(75887)

Write the equation of a line through the points (4,3) and (6,4). Show work.
------------
slope = (4-3)/(6-4) = 1/2
-------------------------------
Form: y = mx + b
Solve for "b":
4 = (1/2)(6) + b
b = 1
----
Equation
y = (1/2)x + 1
===========================
Cheers,
Stan H.